A flywheel with a radius of 0.3 m starts from rest and accelerates with a constant angular acceleragion of 0.60 rad/s square. Calculate the magnitude of the tangential acceleration, radial acceleration and the resultant acceleration of a pount on its rim after it has tur ed 38.96 degree.


1
Expert's answer
2021-12-12T16:40:03-0500

"a_{\tau}=\epsilon r=0.6\cdot 0.3=0.18\ (m/s^2)" .Answer


"a_r=v^2/r=\omega^2r"


"\phi=\omega^2/(2\epsilon)\to \omega^2=2\epsilon \phi"


"a_r=2\epsilon \phi r=2\cdot 0.6\cdot 0.68\cdot 0.3=0.245\ (m/s^2)" .Answer


"a=\sqrt{0.18^2+0.245^2}=0.3\ (m/s^2)"